# Topological transition in cyanobacteria: from motion to structure

### v1

@@ II.1 Model for gliding cyanobacteria
Equations (6) and (7) are discretized in time using the Euler–Maruyama method with a timestep of $\Delta t=10^{-3}$ while equation (5) is discretized as discussed above and shown in Fig. 1.
II.2 Dynamics and transient
Figure 5: Collective behavior of the system at different area coverage $\Phi$ and Péclet numbers $\mathrm{Pe}$. The individual panels show simulation snapshots at $t=150$ in a $30\times 30$ domain with periodic boundary conditions. Filament length is kept constant at $L=1$. The panels’ frames indicate the system’s average nematic order $\langle S\rangle$, calculated as an average of $1\times 1$ sized blocks. With increasing Péclet number $\mathrm{Pe}$ the system transitions from a disordered state into one showing reticulate patterns of nematic bundles.
Figure 6: Average nematic order parameter $\langle S\rangle$ as a function of Péclet number, calculated as an average of $1\times 1$ sized blocks at different area coverages $\Phi$. The filament length is kept constant at $L=1$. With increasing $\Phi$ the transition from disorder to order happens at lower Péclet numbers $\mathrm{Pe}$.
We first investigate the dynamical aspects of the collective self-organization of cyanobacteria filaments. At sufficiently large number density $\rho$ and Péclet number $\mathrm{Pe}$ trichomes self-organize into reticulate patterns [59, 70] as shown in Fig. 2. We study here the process by which this reticulate pattern emerges from an isotropic initial condition, and compare model predictions of the transient dynamics with experimental data. In simulations, we prepare a system of filaments at area coverage $\Phi=1$ with a homogeneous distribution of filaments positions and isotropic distribution of their orientations, see Fig. 2.
@@ II.2 Dynamics and transient
In experiments, Oscillatoria lutea (SAG 1459-3) was cultivated in a medium of BG11 broth diluted 1:100 with deionised water, at $20\pm 1^{\circ}$ with a $10\pm 2$ $\mu$mol m-2 s-1 photon flux on a 16 h day + 8 h night cycle, following Ref. [70]. For imaging, cultures were dispersed by gentle agitation and added dropwise into a 6-well plate (34 mm well diameter), three-quarters filled with medium. Tile-scans of the well were collected every 30 minutes using a confocal laser scanning microscope (Leica TCS SP5) in fluorescence mode (514 nm excitation, 620–780 nm observation). Images of the colony dynamics were cropped and contrast-enhanced to show the central 17$\times$17 mm2 of the well plate, and are given in Fig. 2.
@@
In both experiments and simulations, the aligning interactions lead to the formation of short, dense bundles of filaments. Over time, these bundles connect and self-organize into a network topology. During an initial transient period, the typical separation length between bundles, which provides a characteristic length scale for the patterns, slowly coarsens. After a few hours in the experiments, or simulation times of $t\sim 50$, the pattern then converges into a steady-state configuration of reticulate bundles of high local density, with less dense regions in between. The comparison of the transient dynamics between experiments and simulations reveals that our model quantitatively reproduces the emergence of the reticulate pattern in time; in fact, for $\tau=8$ min, as measured in Ref. [70] a simulated time $t=75$ corresponds to 10 hours, which allows for a direct, one-to-one comparison to the experimental time scale, as in Fig. 2.
@@
$\langle(\Delta\bm{x})^{2}\rangle=\left\langle\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}|\bm{x}_{i}(t_{0}+\Delta t)-\bm{x}_{i}(t_{0})|^{2}\right\rangle\,,$ (8)
@@
where the angular brackets indicate an average over initial times $t_{0}$. The MSD for a polymeric filament can in principle be computed for any part of the filaments, but here, for simplicity, we only use only the positions of the filament heads, $\bm{x}_{i}(t)$. Figure 3(a) shows the dependence of the MSD on the observation interval $\Delta t$. Over short periods there is a ballistic regime, where the MSD grows as $\langle(\Delta\bm{x})^{2}\rangle\propto(\Delta t)^{2}$. On longer timescales, the MSD exhibits a crossover to a diffusive regime, where $\langle(\Delta\bm{x})^{2}\rangle=4D_{\mathrm{eff}}\,\Delta t$, for an effective translational diffusion constant $D_{\mathrm{eff}}$. The inset to Fig. 3(a) shows that our simulations have comfortably reached this diffusive regime, since $\langle(\Delta\bm{x})^{2}\rangle/4\Delta t\to\mathrm{const}$. Figure 3(a) also shows the dependence of the MSD on the Péclet number $\mathrm{Pe}$. As $\mathrm{Pe}$ increases, the crossover time between the ballistic and diffusive regimes grows as $t_{\times}=4D_{\mathrm{eff}}/v_{0}^{2}$. Furthermore, the effective diffusivity $D_{\mathrm{eff}}$ grows with $\mathrm{Pe}$, as detailed in Fig. 3(b). To a good approximation we find that $D_{\mathrm{eff}}$ increases quadratically with $\mathrm{Pe}$, and more specifically as $D_{\mathrm{eff}}=\frac{1}{2}v_{0}^{2}\mathrm{Pe}^{2}$, shown as a dashed line in Fig. 3(b).
It is interesting to compare these results with the behavior of 2D active Brownian particles, whose effective diffusivity $D_{\mathrm{eff}}^{\mathrm{ABP}}=D_{T}+\frac{v_{0}^{2}}{2D_{R}}$, where $D_{T}$ and $D_{R}$ are the translational and rotational diffusion coefficients, respectively [85]. From our definitions of the Péclet number and $\delta\kappa$ it follows that $\mathrm{Pe}=(D_{\omega}\tau^{3})^{-1/2}$. Setting $D_{T}\simeq 0$ and identifying $D_{R}\simeq D_{\omega}\tau^{3}$, we can identify that $D_{\mathrm{eff}}=\frac{1}{2}v_{0}^{2}\mathrm{Pe}^{2}$. This prediction is in good agreement with the simulation results. Figure 3(b) also shows that switching off the interactions among filaments, i.e. setting $g=0$ (which removes the possibility to form bundles), does not affect their translational diffusivity. In other words, the presence of the reticulate pattern does not impede the motility of the filaments; that implies that the filaments do not remain confined within any particular bundle, loop, or other local feature. We note that these results deviate from the prediction of $D_{\mathrm{eff}}\propto\mathrm{Pe}^{2}$ at small values of $\mathrm{Pe}$. This is due to the large values of $\delta\kappa$, that is, the filaments experience a broad range of rotational diffusivities. It is also worth mentioning that our model does not predict any dependence of the MSD on the length $L$ of the filaments, as shown in Fig. 3(c).
@@ II.3 Structure
Figure 7: Collective behavior of the system for different filament lengths $L$ and Péclet numbers $\mathrm{Pe}$. The individual panels show simulation snapshots at $t=150$ in a $30\times 30$ domain with periodic boundary conditions. The density of simulations is set to ensure an equal area coverage, such that $\Phi=1$. The panels’ frames indicate the system’s average nematic order $\langle S\rangle$, calculated as an average of $1\times 1$ sized blocks. With increasing Péclet number $\mathrm{Pe}$ the system transitions from a disordered state into one showing reticulate patterns of nematic bundles. The length scale of the emerging patterns is not affected by filament length.
@@
We now turn to the investigation of the structural properties of the reticulate pattern. Given the nematic nature of the bundles (Fig. 4), and the fact that they can emerge from an isotropic distribution of filament orientations (Fig. 2), we quantify the topological transition to the reticulate network topology using the 2D nematic order parameter $\langle S\rangle=\langle\cos(2\hat{\theta})\rangle$. To calculate $\langle S\rangle$, we sample the local orientation $\hat{\theta}$ of every filament at $160L$ regularly spaced points along the filaments’ length $L$, and calculate $\langle S\rangle$ as a block average within blocks of size $1\times 1$. Here, $\hat{\theta}$ is now the filaments’ local orientation, relative to the average orientation, or director, in each block. At this scale, the filament density is relatively homogeneous, but the blocks are large enough to provide good statistics.
@@
Figure 5 shows snapshots of the steady state of simulations at different Péclet number $\mathrm{Pe}$ and area coverage $\Phi$. The average nematic order $\langle S\rangle$ is given as a colored frame around these snapshots. At constant $\Phi$, as $\mathrm{Pe}$ increases, the system shows a continuous transition from an isotropic disordered state to one of relatively high nematic order, driven first by the formation of bundles, and then by the emergent network-like topology (see Fig. 6). Upon increasing $\Phi$, the transition to a reticulate pattern occurs at lower values of $\mathrm{Pe}$, on account of the higher probability of the filaments to interact. We have previously shown [70] that the typical length scale of the reticulate pattern is directly connected to the curvature fluctuations; in our rescaled units, this predicted typical length $\ell_{\mathrm{pred}}=(\delta\kappa)^{-1}\propto\mathrm{Pe}$. Thus, as $\mathrm{Pe}$ increases, the characteristic size of the reticulate features increases.
@@
Figure 8: Average nematic order parameter $\langle S\rangle$ as a function of Péclet number, calculated as a block-average using $1\times 1$ sized blocks at different filament lengths $L$. The number of filaments is adjusted to ensure an equal area coverage of $\Phi=1$, across all panels. With decreasing $L$, the transition from disorder to order takes on the appearance of a continuous transition.
@@
Figure 9: Filament head-head pair correlation functions (PCFs) parallel to the direction of motion, $g_{\parallel}(r)$ (dashed lines), and orthogonal to it, $g_{\perp}(r)$ (solid lines), see Eqs. (9)-(10) quantify the structure of the reticulate pattern. (a) The time evolution (color axis) of these functions is shown for a system with area coverage $\Phi=1$, filament length $L=1$ and Péclet number $\mathrm{Pe}=3.5$. As time advances and the reticulate pattern forms, $g_{\parallel}(r)$ and $g_{\perp}(r)$ converge to a steady state. For a system with a distribution of curvatures $\delta\kappa$, the minimum of $g_{\perp}(r)$ coincides with $\mathrm{Pe}=(v_{0}\tau\delta\kappa)^{-1}$, indicated by an arrow on the $x$-axis. (b) Steady-state PCFs are shown for different Péclet numbers $\mathrm{Pe}$, keeping filament length $L=1$ and $\Phi=1$. The minima in $g_{\perp}(r)$ coincide with the length scales predicted from the distributions of curvatures, indicated by color-coded arrows. The shoulder in $g_{\parallel}(r)$ moves to larger distances with increasing $\mathrm{Pe}$ (except for $\mathrm{Pe}=9.9$, where finite size effects are visible). (c) Effect of filament length $L$ (color bar) on the PCFs at $\mathrm{Pe}=3.5$. The number density $\rho$ is scaled here, to fix $\Phi=1$. Independent of $L$, the minimum in $g_{\perp}(r)$ coincides with the system’s Péclet number, as denoted by an arrow. The magnitude of the extrema in $g_{\parallel}$ and $g_{\perp}$ decrease with increasing filament length, indicating less well-defined patterns. (d) Temporal evolution of the typical length scale $\ell^{*}$ associated to the reticulate pattern, obtained from the minimum of $g_{\perp}(r)$ in panel (a). At long times, $\ell^{*}\to 3.5$, which is the value of $\mathrm{Pe}$. (e) Dependence of $\ell^{*}$ on $\mathrm{Pe}$ showing that the minimum of $g_{\perp}(r)$ matches the Péclet number and is a good measure of the reticulate’s typical length scale; the dashed line corresponds to the identity.
@@
A powerful method to obtain structural information in soft-matter systems is the pair correlation function (PCF) $g(r)$ [86]. This is defined as the probability of finding another particle within a volume $dV$ at distance $r$ from a reference particle. In polymer physics, it is possible to compute the PCF for any pair of monomers within the same chain or between different chains [87]. In our system, a simple choice is to consider the head-head PCF, where only the head bead of each filament is considered. Furthermore, given the long persistence length of the filament it is convenient to consider $g(r)$ along the directions parallel and normal to the direction of motion of the filament’s head. We thus define
@@
$g_{\parallel}(r)=$ $\frac{1}{2\rho N_{t}w\Delta r}\left\langle\sum_{i,j}\delta\left(|(\bm{x}_{i}-\bm{x}_{j})\cdot\bm{t}_{i}^{0}|-r\right)\right\rangle\,,$ (9)
@@
$g_{\perp}(r)=$ $\frac{1}{2\rho N_{t}w\Delta r}\left\langle\sum_{i,j}\delta\left(|(\bm{x}_{i}-\bm{x}_{j})\times\bm{t}_{i}^{0}|-r\right)\right\rangle\,,$ (10)
@@
as the head-head parallel ($\parallel$) and perpendicular ($\perp$) PCF, respectively. In Eq. (9)-(10) $\delta(r)$ is the Dirac delta distribution, the sums extend to all distinct pairs of filaments $i$ and $j$, and $\bm{t}_{i}^{0}$ is the tangent unit vector to the head bead; we sample filaments’ heads within a narrow rectangle of width $w=10\sigma$ either parallel or perpendicular to the reference filament’s head, and discretize the scalar distance $r$ in bins of size $\Delta r$.
@@
Figure 9(a)) shows the emergence of the reticulate structure over time, as quantified by $g_{\parallel}(r)$ and $g_{\perp}(r)$. As the system starts in an isotropic state (see Fig. 2), initially $g_{\parallel}(r)=g_{\perp}(r)=1$. As the reticulate pattern forms, $g_{\parallel}(r)$ develops a pronounced shoulder, and $g_{\perp}(r)$ develops a distinct minimum. The significance of these features can be related to the reticulate geometry. For example, moving forward from any particular filament’s head, along the direction of its tangent vector, one will be more likely to find other heads so long as they are within a bundle of filaments. Once the bundle bends away, $g_{\parallel}(r)$ will decrease. This is reflected in a shoulder in $g_{\parallel}(r)$. Considering instead exploring in the direction normal to the path of a reference filament’s head, one is also more likely to find other heads, but only so long as they are within the same bundle of filaments. As the bundles are relatively narrow, this probability decays quickly and drops below one in the regions between bundles, but then grows back when another bundle is encountered; this arrangement produces a minimum in $g_{\perp}(r)$. The position $\ell^{*}$ of this minimum correlates with the pattern’s length scale; as the pattern coarsens, the depth of the minimum in $g_{\perp}(r)$ and height of the shoulder in $g_{\parallel}(r)$ increase, indicating the growth of the pattern length scale, as was observed in Fig. 2. The position of the minimum eventually settles on a value close to that of $\ell_{\mathrm{pred}}=\mathrm{Pe}$.
@@
Figure 9(d)-(e) show the dependence of $\ell^{*}$ on time and Péclet number, respectively, as based on the the results in Fig. 9(a)-(b). The position of the minimum of $g_{\perp}(r)$ converges to the Péclet number of the simulation, which in our rescaled units, corresponds to the prediction for the length scale of the pattern $\ell_{\mathrm{pred}}=\mathrm{Pe}$ [70].
III Conclusions and Outlook

### v2

@@ II.2 Simulations
II.3 Experiments
@@
Oscillatoria lutea (SAG 1459-3) was cultivated in a medium of BG11 broth diluted 1:100 with deionised water, at $20\pm 1^{\circ}$ with a $10\pm 2$ $\mu$mol m-2 s-1 photon flux on a 16 h day + 8 h night cycle, following Ref. [72]. For imaging, cultures were dispersed by gentle agitation and added dropwise into a 6-well plate (34 mm well diameter), three-quarters filled with medium. Tile-scans of wells were collected using a confocal laser scanning microscope (Leica TCS SP5) in fluorescence mode (514 nm excitation, 620–780 nm observation). Images were processed using ImageJ and MATLAB.
[…]
III Results
III.1 Dynamics and transient
Figure 6: Continuous transition to reticulate pattern. The dependence of the average nematic order parameter $\langle S\rangle$ on Péclet number is calculated as an average of $1\times 1$ sized blocks at different area coverages $\Phi$. The filament length is kept constant at $L=1$. With increasing $\Phi$ the transition from disorder to order happens at lower Péclet numbers $\mathrm{Pe}$. Error bars representing the standard error are smaller than symbol size.
Figure 7: Effect of varying the interaction strength $g$. Snapshots of the steady-state simulated system for varying dimensionless interaction strength $g$ (see legends) at fixed $\mathrm{Pe}=3.5$, $L=1$, and $\Phi=1$. Upon increasing $g$, the colony transitions from an isotropic distribution of filaments (low $g$ limit) to a reticulate pattern ($g\approx 2$), to aster-like configurations ($g\approx 4$), and finally to strong coils ($g\gtrsim 8$).
We first investigate the dynamical aspects of the collective self-organization of cyanobacteria filaments. At sufficiently large number density $\rho$ and Péclet number $\mathrm{Pe}$ trichomes self-organize into reticulate patterns [61, 72] as shown in Fig. 2. We study here the process by which this reticulate pattern emerges from an isotropic initial condition, and compare model predictions of the transient dynamics with experimental data. In simulations, we prepare a system of filaments at area coverage $\Phi=1$ with a homogeneous distribution of filaments positions and isotropic distribution of their orientations, see Fig. 2. For the corresponding experiments, images of the colony dynamics were cropped and contrast-enhanced to show the central 17$\times$17 mm2 of the well plate, and are given in Fig. 2.
@@ III.1 Dynamics and transient
In both experiments and simulations, the aligning interactions lead to the formation of short, dense bundles of filaments. Over time, these bundles connect and self-organize into a network topology. During an initial transient period, the typical separation length between bundles, which provides a characteristic length scale for the patterns, slowly increases. After a few hours in the experiments, or simulation times of $t\sim 50$, the pattern then converges into a steady-state configuration of reticulate bundles of high local density, with less dense regions in between. The comparison of the transient dynamics between experiments and simulations reveals that our model quantitatively reproduces the emergence of the reticulate pattern in time; in fact, for $\tau=8$ min, as measured in Ref. [72] a simulated time $t=75$ corresponds to 10 hours, which allows for a direct, one-to-one comparison to the experimental time scale, as in Fig. 2.
@@
$\langle(\Delta\mathbf{x})^{2}\rangle=\left\langle\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}|\mathbf{x}_{i}(t_{0}+\Delta t)-\mathbf{x}_{i}(t_{0})|^{2}\right\rangle\,,$ (8)
@@
where the angular brackets indicate an average over initial times $t_{0}$. The MSD for a polymeric filament can in principle be computed for any part of the filaments, but here, for simplicity, we only use the positions of the filament heads, $\mathbf{x}_{i}(t)$. Figure 3(a) shows the dependence of the MSD on the observation interval $\Delta t$. Over short periods there is a ballistic regime, where the MSD is dominated by the filaments’ self-propulsion and grows as $\langle(\Delta\mathbf{x})^{2}\rangle\approx v_{0}^{2}(\Delta t)^{2}$. On longer timescales, the MSD exhibits a crossover to a diffusive regime (akin to an active Brownian particle), where $\langle(\Delta\mathbf{x})^{2}\rangle=4D_{\mathrm{eff}}\,\Delta t$, for an effective translational diffusion constant $D_{\mathrm{eff}}$. The inset to Fig. 3(a) shows that our simulations have comfortably reached this diffusive regime, since $\langle(\Delta\mathbf{x})^{2}\rangle/4\Delta t\to\mathrm{const}$. Figure 3(a) also shows the dependence of the MSD on the Péclet number $\mathrm{Pe}$. As $\mathrm{Pe}$ increases, the crossover time between the ballistic and diffusive regimes grows as $t_{\times}=4D_{\mathrm{eff}}/v_{0}^{2}$ (see Fig. 3(a) inset).
The temporal dynamics of the angle are complex. However, because the MSD of an Ornstein–Uhlenbeck process is known [89], we can identify the long-time rotational diffusion constant as $D_{R}\simeq D_{\omega}\tau^{2}$, for $t\gg\tau$. A self-propelled particle in the absence of translational diffusion and with constant speed, whose orientation is characterized by a diffusion constant $D_{R}$ will exhibit a long-time diffusive behavior with an effective diffusion constant $D_{\text{eff}}=\frac{v_{0}^{2}}{2D_{r}}$ (in 2D) [90]. Using these observations, we can estimate that the filaments’ heads will show a long-time diffusive behavior with $D_{\mathrm{eff}}=\frac{1}{2}v_{0}^{2}\tau\mathrm{Pe}^{2}$. This prediction is in good agreement with the simulation results, shown as a dashed line in Fig. 3(b), for large Péclet numbers. At low $\mathrm{Pe}$, we observe some deviation from the predicted scaling, due to the fact that the estimated $D_{R}$ emerges from the long-time average of the curvature fluctuations quantified by $D_{\omega}$. At low $\mathrm{Pe}$, we have strong curvature fluctuations ($\delta\kappa\sim\mathrm{Pe}^{-1}$), thus the orientational dynamics cannot be approximated as a diffusive process, because the correlation time $\tau$ cannot be ignored in this limit.
Figure 3(b) also shows that being ordered into bundles has a small but noticeable effect on their mobility, since switching off the interactions among filaments, i.e. setting $g=0$ (which removes the possibility to form bundles) causes a small increase in their effective long-time translational diffusion $D_{\text{eff}}$. In other words, the presence of the reticulate pattern alters the motility of the filaments in a subtle way but without disrupting it; that implies that the filaments do not remain confined within any particular bundle, or loop.
It is also worth mentioning that our model does not predict any dependence of the MSD on the length $L$ of the filaments, as shown in Fig. 3(c).
@@ III.2 Structure
Figure 8: Emergence of reticulate patterns. The mosaic plot shows the steady-state collective behavior of the system for different filament lengths $L$ and Péclet numbers $\mathrm{Pe}$. The individual panels show simulation snapshots at steady state in a $30\times 30$ domain with periodic boundary conditions. The density of simulations is set to ensure an equal area coverage, such that $\Phi=1$. The panels’ frames indicate the system’s average nematic order $\langle S\rangle$, calculated as an average of $1\times 1$ sized blocks. With increasing $\mathrm{Pe}$, the system transitions from a disordered state into one showing reticulate patterns of nematic bundles. The length scale of the emerging patterns is not affected by filament length.
@@
We now turn to the investigation of the structural properties of the reticulate pattern. Given the nematic nature of the bundles (Fig. 4), and the fact that they can emerge from an isotropic distribution of filament orientations (Fig. 2), we quantify the topological transition to the reticulate network topology using the 2D nematic order parameter $\langle S\rangle$. To calculate $\langle S\rangle$, we sample the local orientation $\hat{\mathbf{t}}_{i}$ of every filament at $160L$ regularly spaced points along the filaments’ length $L$, and calculate the local nematic tensor $\mathbb{Q}=\frac{1}{n}\left(\sum_{i}2\mathbf{t}_{i}\otimes\mathbf{t}_{i}-\mathbb{I}\right)$ within blocks of size $1\times 1$, where $n$ is the number of samples within a block. At this scale, the filament density is relatively homogeneous, but the blocks are large enough to provide good statistics. The local order parameter $S_{\text{loc}}$ is the largest eigenvalue of $\mathbb{Q}$. $\langle S\rangle$ is an average over all blocks. Local nematic order $S_{\text{loc}}$ is high within bundles and lower in the dilute regions between them (see Fig. 5 inset at $\mathrm{Pe}=3.5$ and $\Phi=1$).
@@
Figure 5 shows snapshots of the steady state of simulations at different Péclet number $\mathrm{Pe}$ and area coverage $\Phi$. The average nematic order $\langle S\rangle$ is given as a colored frame around these snapshots. The transition from disorder to order of our cyanobacteria colony as density increases has been explored previously [72]. Here, we observe a transition also dependent on the Péclet number. At constant $\Phi$, as $\mathrm{Pe}$ increases, the system shows a continuous transition from an isotropic disordered state to one of relatively high nematic order, driven first by the formation of bundles, and then by the emergent network-like topology (see Fig. 6). Upon increasing $\Phi$, the transition to a reticulate pattern occurs at lower values of $\mathrm{Pe}$, on account of the higher probability of the filaments to interact. We have previously shown [72] that the typical length scale of the reticulate pattern is directly connected to the curvature fluctuations; in our rescaled units, this predicted typical length $\ell_{\mathrm{pred}}=(\delta\kappa)^{-1}=\mathrm{Pe}\,v_{0}\tau$. Thus, as $\mathrm{Pe}$ increases, the characteristic size of the reticulate features increases.
The second dimensionless parameter governing the filaments’ dynamics is the interaction strength $g$. We explore how varying $g$ affects the structure formation in Fig. 7. In the limit of $g\to 0$, the system remains isotropic. Upon increasing $g$, the reticulate pattern emerges ($g\approx 2$). This is the parameter range that most closely resembles the experimental observations of young O. lutea colonies. For larger values of $g$, $g\approx 4$, the interactions dominate over the curvature fluctuations, and we observe aster-like structures (see e.g.  [91, 92], though in a different physical context), and strong coils for $g\gtrsim 8$ (see e.g. [93]).
@@
Figure 9: Approach to the continuous transition. The dependence of the average nematic order parameter $\langle S\rangle$ on Péclet number is calculated as a block-average using $1\times 1$ sized blocks at different filament lengths $L$. The number of filaments is adjusted to ensure an equal area coverage of $\Phi=1$, across all panels. With decreasing $L$, the transition from disorder to order takes on the appearance of a continuous transition.
@@
Figure 10: Emergence of structure. Filament head-head pair correlation functions (PCFs) parallel to the direction of motion, $g_{\parallel}(r)$ (dashed lines), and orthogonal to it, $g_{\perp}(r)$ (solid lines), see Eqs. (9)-(10), quantify the structure of the reticulate pattern. (a) The time evolution (color axis) of these functions is shown for a system with area coverage $\Phi=1$, filament length $L=1$ and Péclet number $\mathrm{Pe}=3.5$. As time advances and the reticulate pattern forms, $g_{\parallel}(r)$ and $g_{\perp}(r)$ converge to a steady state. For a system with a distribution of curvatures $\delta\kappa$, the minimum of $g_{\perp}(r)$ coincides with $\mathrm{Pe}=(v_{0}\tau\delta\kappa)^{-1}$, indicated by an arrow on the $x$-axis. (b) Steady-state PCFs are shown for different $\mathrm{Pe}$, keeping filament length $L=1$ and $\Phi=1$. The minima in $g_{\perp}(r)$ coincide with the length scales predicted from the distributions of curvatures, indicated by color-coded arrows. The shoulder in $g_{\parallel}(r)$ moves to larger distances with increasing $\mathrm{Pe}$ (except for $\mathrm{Pe}=9.9$, where finite size effects are visible). (c) Effect of filament length $L$ (color bar) on the PCFs at $\mathrm{Pe}=3.5$. The number density $\rho$ is scaled here, to fix $\Phi=1$. Independent of $L$, the minimum in $g_{\perp}(r)$ coincides with the system’s $Pe$, as denoted by an arrow. The magnitude of the extrema in $g_{\parallel}$ and $g_{\perp}$ decrease with increasing filament length, indicating less well-defined patterns. (d) Temporal evolution of the typical length scale $\ell^{*}$ associated to the reticulate pattern, obtained from the minimum of $g_{\perp}(r)$ in panel (a). At long times, $\ell^{*}\to 3.5$, which is the value of $\mathrm{Pe}$. (e) Dependence of $\ell^{*}$ on $\mathrm{Pe}$ showing that the minimum of $g_{\perp}(r)$ matches the Péclet number and is a good measure of the reticulate’s typical length scale; the dashed line corresponds to the identity.
@@
A powerful method to obtain structural information in soft-matter systems is the pair correlation function (PCF) $g(r)$ [94]. This is defined as the probability of finding another particle within a volume $dV$ at distance $r$ from a reference particle. In polymer physics, it is possible to compute the PCF for any pair of monomers within the same chain or between different chains [95]. In our system, a simple choice is to consider the head-head PCF, where only the head bead of each filament is considered. Furthermore, given the long persistence length of the filament it is convenient to consider $g(r)$ along the directions parallel and normal to the direction of motion of the filament’s head. We thus define the head-head parallel PCF
@@
$g_{\parallel}(r)=$ $\frac{1}{2\rho N_{t}w\Delta r}\left\langle\sum_{i,j}\delta\left(|(\mathbf{x}_{i}-\mathbf{x}_{j})\cdot\mathbf{t}_{i}^{0}|-r\right)\right\rangle\,,$ (9)
and the head-head perpendicular PCF
@@
$g_{\perp}(r)=$ $\frac{1}{2\rho N_{t}w\Delta r}\left\langle\sum_{i,j}\delta\left(|(\mathbf{x}_{i}-\mathbf{x}_{j})\times\mathbf{t}_{i}^{0}|-r\right)\right\rangle\,.$ (10)
@@
In Eq. (9)-(10) $\delta(r)$ is the Dirac delta distribution, the sums extend to all distinct pairs of filaments $i$ and $j$, and $\mathbf{t}_{i}^{0}$ is the tangent unit vector to the head bead; we sample filaments’ heads within a narrow rectangle of width $w=10\sigma$ either parallel or perpendicular to the reference filament head’s orientation, and discretize the scalar distance $r$ in bins of size $\Delta r$.
@@
Figure 10(a)) shows the emergence of the reticulate structure over time, as quantified by $g_{\parallel}(r)$ and $g_{\perp}(r)$. As the system starts in an isotropic state (see Fig. 2), initially $g_{\parallel}(r)=g_{\perp}(r)=1$. As the reticulate pattern forms, $g_{\parallel}(r)$ develops a pronounced shoulder, and $g_{\perp}(r)$ develops a distinct minimum. The significance of these features can be related to the reticulate geometry. For example, moving forward from any particular filament’s head, along the direction of its tangent vector, one will be more likely to find other heads so long as they are within a bundle of filaments. Once the bundle bends away, $g_{\parallel}(r)$ will decrease. This is reflected in a shoulder in $g_{\parallel}(r)$. Considering instead exploring in the direction normal to the path of a reference filament’s head, one is also more likely to find other heads, but only so long as they are within the same bundle of filaments. As the bundles are relatively narrow, this probability decays quickly and drops below one in the regions between bundles, but then grows back when another bundle is encountered; this arrangement produces a minimum in $g_{\perp}(r)$. The position of this minimum $\ell^{*}=\operatorname{arg\,min}g_{\perp}(r)$ serves as a measurement of the pattern’s length scale; as the pattern coarsens, the depth of the minimum in $g_{\perp}(r)$ and height of the shoulder in $g_{\parallel}(r)$ increase, indicating the growth of the pattern length scale, as was observed in Fig. 2. The position of the minimum eventually settles on a value close to that of $\ell_{\mathrm{pred}}=\mathrm{Pe}\,v_{0}\tau$.
@@
We stress that, despite the fact that our system is topologically non-trivial and thus violates the condition of homogeneity used to derive the PCF [96], the latter does reflect important structural information about the filament colony.
@@
Figure 10(d)-(e) show the dependence of $\ell^{*}$ on time and Péclet number, respectively, as based on the results in Fig. 10(a)-(b). The position of the minimum of $g_{\perp}(r)$ converges to the Péclet number of the simulation, which in our rescaled units, corresponds to the prediction for the length scale of the pattern $\ell_{\mathrm{pred}}=\mathrm{Pe}\,v_{0}\tau$ [72].
IV Conclusions
We have studied the dynamics and structural properties of a colony of filamentous cyanobacteria as they undergo a topological transition to a reticulate pattern. Their ability to glide over one another while experiencing weak aligning interactions enables the formation of a disordered network-like structure of filaments within bundles that link together. In addition to the network topology, confined cyanobacteria systems can form topological defects [76, 75], bearing a striking resemblance to some active nematic systems [97, 98]. Other than these confined systems, where defects can only resolve the mismatch of order by non-local reconfiguration, the unconfined cyanobacteria can glide over one another, resulting in dynamics not dominated by topological defects.